OpenStudy (anonymous):

evaluate the function as indicated. Determine its domain and range. x^2+2, x<1. 2x^2+2, x>1. A: f(-2). B: f(0). C: f(1). D: f(s^2+2).

OpenStudy (anonymous):

@ash2326 @waterineyes I only need help basically with D.

OpenStudy (anonymous):

A,B,C I plugged those values into the first function..since they are either less than or equal to 1!

OpenStudy (anonymous):

There are two functions given.. What we have to do first I am not getting... Second part of your question is clear I am not getting the very first statement..

OpenStudy (anonymous):

I have to plug in the values, lets say f(-2) into the appropriate function first.... for a,b,c, and d. lastly, i find their domain and range..

OpenStudy (anonymous):

since -2 is less than 1, i would plug that into the first function.. ok?

OpenStudy (anonymous):

Yes you are going right.. Go ahead.. by putting x = -2 what you got??

OpenStudy (anonymous):

i got 6, for f(-2).. I'm not sure thats right, but i am pretty sure

OpenStudy (anonymous):

Have faith on yourself dear.. You are right.. Now do the same by putting x = 0..

OpenStudy (anonymous):

Ok. for f(0) i got 2

OpenStudy (anonymous):

Yes this one is also correct.. Now find f(1) by putting x = 1.. But be careful do you know where it is to be plugged in?

OpenStudy (anonymous):

first one, since it is also equal to.. i got 3

OpenStudy (anonymous):

But in the question either < is given or > is given.. I did not see any equal sign there with inequality..

OpenStudy (anonymous):

I forgot tp include, sorry. equal sign should be in first function only

OpenStudy (anonymous):

with that being said @waterineyes , can u help with D?

OpenStudy (ash2326):

@schmidtdancer I'll help you with D

OpenStudy (anonymous):

Then you are right..

OpenStudy (anonymous):

kk thx guys

OpenStudy (ash2326):

s^2+2, s^2 is always positive so, s^2+2>2 for all s values so just put x=s^2+2 in the f(x) given for x>1

OpenStudy (anonymous):

kk so plug that into the second function

OpenStudy (anonymous):

See as you have plugged for x= 1, -2 etc etc now replace x by \(s^2 + 2\)..

OpenStudy (ash2326):

Yeah \[f(s^2+2)=(s^2+2)^2+2\]

OpenStudy (anonymous):

Yes you have to plug it in second because \(s^2 + 2 > 0\)..

OpenStudy (anonymous):

yay thats what i had!!

OpenStudy (anonymous):

Show your work..

OpenStudy (anonymous):

shouldn't there be a 2 in front of that @ash2326

OpenStudy (anonymous):

like 2(s^2+2)^2 + 2

OpenStudy (ash2326):

oh I missed the 2 \[f(s^2+2)=2(s^2+2)^2+2\]

OpenStudy (anonymous):

Yes you are right..

OpenStudy (anonymous):

im not sure how to simplify the rest

OpenStudy (anonymous):

For simplification you can use the formula: \[(a+b)^2 = a^2 + b^2 + 2ab\]

OpenStudy (anonymous):

2s^4+6..thats what I got when I did that but I honestly have no idea

OpenStudy (anonymous):

\(a = s^2\) and \(b = 2\) there.. Just use the formula and try to do at least this step only..

OpenStudy (anonymous):

s^4+4 +2ab?

OpenStudy (anonymous):

Why you did replace a and b in 2ab term??

OpenStudy (anonymous):

Sorry did not..

OpenStudy (anonymous):

s^4+4+2(s^2)(2)

OpenStudy (anonymous):

Yes... Now put this in that expression..

OpenStudy (anonymous):

4s^2+s^4+4

OpenStudy (anonymous):

Yes now put it in original expression we have just found..

OpenStudy (anonymous):

which is what agin?

OpenStudy (anonymous):

\[= 2(s^4 + 4s^2 + 4) + 2\] can you solve this ??

OpenStudy (anonymous):

that is 2s^4+8s^2+4

OpenStudy (anonymous):

?

OpenStudy (anonymous):

you are confusing me now... Wait.. Just cool your mind for 30 seconds...

OpenStudy (anonymous):

Wait, I thought that was right?

OpenStudy (anonymous):

2(s4+4s2+4)+2 <-- this equals what I put

OpenStudy (anonymous):

Yes now you got it.. Now solve it and tell me what you get..

OpenStudy (anonymous):

I got 2s^4+8s^2+4....

OpenStudy (anonymous):

+10 i mean

OpenStudy (anonymous):

Yes now are right.. Finally you did that..

OpenStudy (anonymous):

So 2s^4+8s^2+10 is the answer

OpenStudy (anonymous):

yes this is value of \(f(s^2 + 2)\)..